# Inference-Time Text-to-Video Alignment with Diffusion Latent Beam Search

### v1

@@ 5.4 Evaluation with AI Feedback
As discussed in Section 4, we obtain a manageable reward function through the reward calibration, which reduces the cost for frequent evaluation queries in inference-time search. While DLBS efficiently improves the calibrated reward, (Figure 5), a natural question is whether DLBS can also improve the original VLM feedback by optimizing their calibrated rewards. We use each calibrated reward (Figure 4) for DLBS and evaluate the quality with Gemini or GPT-4o. Figure 9 demonstrates that DLBS maximizing calibrated rewards can improve the original preference feedback from VLMs, as we grow the search budget. This highlights that our recipe to design inference-time search, combining reward calibration and beam search, can contribute to improving video quality in general.
Figure 9: VLM evaluation of video generated with DLBS. As in Figure 5, we use each calibrated reward (Figure 4) for DLBS and evaluate the quality with Gemini or GPT-4o. DLBS on calibrated reward can improve the original preference feedback from VLMs.
5.5 Alignment–Diversity Tradeoff
Alignment for diffusion models can steer desirable outputs, but it is said that the diversity of generated samples or the performance of original models often degrade (Lee et al., 2023b; Wu et al., 2024b). While inference-time search does not change or degrade the model itself, we here compare the diversity of samples among BoN, GS, and DLBS. We measure the sample diversity as a mean pairwise distance of ViCLIP embeddings (see Appendix J). Figure 10 reveals that DLBS achieves high performance with higher diversity than BoN or GS. This exhibits a benefit from the wider possession and exploration of diffusion latents in DLBS.
@@ 6 Related Works
Classifier guidance (Dhariwal & Nichol, 2021; Ho & Salimans, 2022) has been most popular to enhance text-content alignment. On top of that, recent works (Huang et al., 2024a; Li et al., 2024b) leverage reward or external feedback at inference time by selecting better latents (Wallace et al., 2023b), that probably achieve higher reward, during the reverse process. Kim et al. (2025) propose twisted SMC (Wu et al., 2024b) with reward gradient, which is not suitable for non-differentiable feedback and for the domain such as video where reward gradient needs a huge memory cost. Gradient-free methods (Zheng et al., 2024; Ma et al., 2025) such as SMC (Singhal et al., 2025) or greedy search (Kim et al., 2024) often exhibit sub-optimal results affected by inaccurate reward estimate from noisy latents. Yeh et al. (2024) use ODE to estimate the reward, but it highly depends on Karras sampler (Karras et al., 2022) to avoid numerical instability. In contrast, we address the error propagation from inaccurate reward estimate with beam search and lookahead estimator via deterministic DDIM, which is more popular and stable. Our methods work more scalably when allocating more computation budget at inference time. See Appendix I for further related works.
@@ 7 Discussion and Limitation
We here assume that VLMs work as a proxy of human evaluation, and we demonstrate both qualitative and quantitative improvement of the video quality. It would be important to extend us to scalable methods driven by human feedback not limited to VLMs. While spending more computation at inference time significantly improves perceptual quality, the generation definitely requires more time. It is orthogonal compared to speeding up the sampling process via distillation (Meng et al., 2022; Song et al., 2023), architecture changes (Oshima et al., 2024), or parallel sampling (Shih et al., 2023). We believe both high-quality and speedy sampling have practical needs and should be balanced.
@@ 8 Conclusion
This paper studies which metrics we should optimize and how we can optimize them for better text-to-video generation. We point out that feedback from capable VLMs reflects multiple dimensions of video quality so it is insufficient to optimize an existing metric alone, rather we should calibrate the reward by combination. Our DLBS with LA estimator reduces the error propagation from the inaccurate reward estimate. We demonstrate DLBS is the most scalable and robust inference-time search that significantly improves video quality as we increase the search budget, lookahead, and denoising steps. We hope our work encourages more uses of inference-time computation for text-to-video models.
@@ Acknowledgements
We thank Po-Hung Yeh, Shohei Taniguchi, Kuang-Huei Lee, Arnaud Doucet, Heiga Zen, and Yusuke Iwasawa for their support and helpful discussion on the initial idea of this work. HF was supported by JSPS KAKENHI Grant Number JP22J21582, and MS was supported by JSPS KAKENHI Grant Number J23H04974.
Impact Statement
@@
On the other hand, the ability to generate highly realistic videos raises concerns about the potential for misuse in creating deceptive content, including deepfakes and misinformation. Like other generative models such as large language models, text-to-video models and their inference-time search may inherit and amplify biases present in the training data due to the misalignment. This might lead to the generation of videos that perpetuate harmful stereotypes or underrepresent certain groups.
@@
Appendix A Implementation Details
We use the pre-trained T2V diffusion model Latte (Ma et al., 2024), a T5-conditioned (Raffel et al., 2020) Latent Diffusion Transformer-based framework with 1.1B parameters. Our implementation is based on the publicly available code repository at https://github.com/Vchitect/Latte, and we utilize the released checkpoint from https://huggingface.co/maxin-cn/Latte-1.
Following the original configuration of Latte, we employ a linear noise schedule with $\beta_{\text{start}}=0.0001$ and $\beta_{\text{end}}=0.02$, and set the guidance scale $w_{\text{cfg}}=7.5$ for all experiments.
All experiments are conducted in half precision (fp16) on a single NVIDIA A100 GPU (40 GB) with a batch size of 1. We report generation times for various configurations of the diffusion steps $T$, the lookahead steps $T^{\prime}$, and the search budget $KB$. The approximate time consumption for different settings is summarized in Table 1.
For AI feedback, we use API endpoints: gpt-4o-2024-11-20 and gemini-1.5-pro-002.
Table 1: Approximate video generation time (minutes) for different settings of the diffusion steps $T$, the lookahead steps $T^{\prime}$, and the search budget $KB$ on a single NVIDIA A100 GPU (40 GB) with a batch size of 1, using the Gemini Calibrated Reward for MSRVTT-test.
$T$ $T^{\prime}$ $KB$ Execution Time (min)
50 1 4 11
50 1 8 22
50 1 16 45
50 1 32 90
50 6 8 30
100 1 8 45
200 1 8 90
Appendix B Details of Metric Rewards
Subject Consistency  We adopt the subject consistency metric proposed in VBench (Huang et al., 2024b) to quantify how consistently a subject is depicted across consecutive video frames. Concretely, for each frame $i$ in a video, we extract a feature representation $\mathbf{d}_{i}$ using DINO (Caron et al., 2021) with a ViT-B/16 (Dosovitskiy et al., 2020) backbone. Let $\langle\mathbf{d}_{i},\mathbf{d}_{j}\rangle$ denote the cosine similarity between the features $\mathbf{d}_{i}$ and $\mathbf{d}_{j}$. Then, VBench defines the subject consistency metric as follows:
@@ Appendix B Details of Metric Rewards
Motion Smoothness  We adopt the frame-interpolation-based metric originally proposed in VBench (Huang et al., 2024b) to assess whether a generated video’s motion is smooth and physically plausible. In particular, this metric leverages the motion prior from AMT (Li et al., 2023), employing its AMT-S variant for frame reconstruction. Concretely, let $\bigl{[}\mathbf{f}_{0},\mathbf{f}_{1},\mathbf{f}_{2},\ldots,\mathbf{f}_{2n}\bigr{]}$ denote the frames of a generated video. We remove each odd-numbered frame to obtain a lower-frame-rate sequence $\bigl{[}\mathbf{f}_{0},\mathbf{f}_{2},\mathbf{f}_{4},\ldots,\mathbf{f}_{2n}\bigr{]}$, and rely on AMT-S to reconstruct the missing frames $\bigl{[}\hat{\mathbf{f}}_{1},\hat{\mathbf{f}}_{3},\ldots,\hat{\mathbf{f}}_{2n-1}\bigr{]}$. We then compute the Mean Absolute Error (MAE) between these reconstructed frames and the original odd-numbered frames, denoting this measure by $R_{\text{smoothness}}$. Finally, following the normalization scheme introduced in VBench, we define:
@@
Dynamic Degree  This measure quantifies the overall magnitude of dynamic object movement. Let $T$ be the total number of frames in the generated video. For each pair of consecutive frames $t$ and $t+1$, we estimate the optical flow $\mathbf{v}_{t}$ using RAFT (Teed & Deng, 2020), compute its norm $\|\mathbf{v}_{t}\|$, and sum these values across all frames:
@@
Aesthetic Quality  This criterion evaluates compositional rules, color harmony, and overall artistic merit on a per-frame basis. Concretely, for each frame $i$ in a video, we extract a CLIP image embedding $\mathbf{c}^{image}_{i}$ using the CLIP ViT-L/14 model (Radford et al., 2021). We then feed $\mathbf{c}^{image}_{i}$ into the LAION aesthetic predictor (LAION-AI, 2022), which assigns a raw rating $r_{i}\in[0,10]$. To normalize these scores to the $[0,1]$ range, we set
@@
Imaging Quality  This indicator assesses low-level distortions (e.g., over-exposure, noise, blur) in each generated frame. We adopt the MUSIQ predictor (Ke et al., 2021), trained on the SPAQ dataset (Fang et al., 2020). The frame-wise score is normalized to [0, 1] by dividing by 100, and the final video score is the mean of these normalized values across all frames in the same way as Equation 16 and Equation 17.
Text-Video Consistency This measure captures how closely a generated video’s content aligns with its text prompt. We employ ViCLIP (Xu et al., 2021), a model pre-trained on a 10M video-text dataset and fine-tuned to handle temporal relationships, to embed both the video frames and the text. Since ViCLIP computes embeddings from 8-frame inputs, we sample 8 frames from 16-frame video. Let $\mathbf{v}^{\text{video}}$ denote the resulting video embedding and $\mathbf{v}^{\text{text}}$ denote the text embedding. We then define the final alignment score as the cosine similarity between these embeddings:
@@
Appendix C List of Prompts

### v2

@@ 5.2 Evaluation with AI and Human Feedback
As discussed in Section 4, we obtain a manageable reward function through the reward calibration, which reduces the cost for frequent evaluation queries in inference-time search. While DLBS efficiently improves the calibrated reward (Figure 5), a natural question is whether DLBS can improve an actual assessment by VLMs or humans by optimizing their calibrated rewards. We first use each calibrated reward for DLBS, then evaluate the quality using discrete scores (from 1 to 10) from Gemini or GPT-4o. Figure 7 (Left) demonstrates that DLBS maximizing calibrated rewards can improve the original preference feedback from VLMs, as we grow the search budget.
Next, we conduct an evaluation with VideoScore [54], a metric trained on human judgments that evaluates videos across five quality criteria. Figure 7 (Middle) shows that DLBS significantly improves the quantitative evaluation based on human evaluation. Lastly, we perform pairwise comparisons between DLBS-LA and GS ($KB=1$) by three human evaluators (Figure 7; Right). The results confirm that, whatever models or prompts we choose, the quality of content generated by DLBS-LA consistently outperforms that of a baseline. This emphasizes that our proposed method, integrating reward calibration and beam search in a latent space, effectively enhances perceptual video quality. See Appendix L.10 for the details.
Figure 7: (Left) DLBS on calibrated reward can improve the original preference feedback from VLMs. As in Figure 5, we use each calibrated reward (Figure 4) for DLBS and evaluate the quality with Gemini or GPT-4o. (Middle) DLBS on calibrated reward also improves another qualitative metric, the most, VideoScore [54], which is not involved in a reward calibration. (Right) The pairwise comparisons between DLBS-LA and GS ($KB=1$) by three human raters, when DLBS-LA searches for the best latents in a diffusion process.
5.3 Scaling Model Parameters and Capabilities
We scale up the base diffusion model from Latte to the latest SoTA models such as CogVideoX-5B [19] and Wan 2.1-14B [20], and evaluate if DLBS can improve the generation quality from those larger models. Since base models are more capable here, we adopt more challenging prompts from DEVIL-very-high (22 prompts) and Movie Gen Video Bench [55] (20 prompts). See Appendix F and I.2 for the details. Figure 6 (Right) shows that our methods achieve significant improvement in calibrated reward for both models. As shown in Figure 7 (Right), human evaluation also supports that our proposed method could generally work well with any text-to-video models, even with more capable models in the future. See Appendix L.9 for other results with CogVideoX and Wan 2.1.
Figure 8: (Left) Comparison among different LA steps $T^{\prime}$ on MSRVTT-test ($KB=8$). The performance improves as the number of LA steps increases. (Right) Perceptual quality comparison on a large and complex prompt set (1,003 prompts from Movie Gen Video Bench [55]). DLBS-LA can generalize to diverse prompts.
5.4 Ablation Study
Lookahead Steps for Reward Estimate  We scale up LA steps $T^{\prime}$ to obtain an accurate reward estimate. We use MSRVTT-test ($KB=8$) and Gemini reward for experiments. Figure 8 (Left) shows that as the number of LA steps increases, the performance improves more. Even $T^{\prime}=2,3$ significantly outperforms the posterior mean, which is often used in prior works [15, 11, 41]. This is because the sub-optimal performance of inference-time search comes from the approximation errors and LA estimator can notably reduce them. As shown in Figure 5, DLBS-LA $(KB=8,T^{\prime}=6)$ achieves comparable or even outperforming results with DLBS $(KB=32)$. It is quite beneficial to spend a computation to estimate the reward accurately. See Appendix L.4 for further discussions. In addition, an ablation study on diffusion steps is shown in Appendix L.5.
A Large and Complex Prompt Set  To confirm the robustness of our approach on a large and highly diverse prompt set, we compare GS ($KB{=}1$) with DLBS-LA ($KB{=}8,\;T^{\prime}{=}6$) with all the prompts in Movie Gen Video Bench [55], which comprises 1,003 prompts (Figure 8; Right). We observe that DLBS-LA consistently delivered substantially higher alignment rewards, demonstrating that DLBS generalizes effectively to complex prompt distributions.
@@ 6 Related Works
Classifier guidance [35, 56] has been the most popular to enhance text-content alignment. On top of that, recent works [11, 39] leverage reward or external feedback at inference time by selecting better latents [57], which probably achieve higher rewards during the reverse process. Kim et al. [58] propose twisted SMC [37] with reward gradient, which is not suitable for non-differentiable feedback and for domains such as video, where reward gradient needs a huge memory cost. Gradient-free methods [59, 17] such as SMC [41] or greedy search [15] often exhibit sub-optimal results affected by inaccurate reward estimates from noisy latents. Yeh et al. [60] uses ODE to estimate the reward, but it highly depends on Karras sampler [28] to avoid numerical instability. In contrast, we address the error propagation from inaccurate reward estimates with beam search and lookahead estimator via deterministic DDIM, which is more popular and stable. Our methods work more scalably when allocating more computation budget at inference time. See Appendix M for further related works.
@@ 7 Discussion and Limitation
Our reward calibration assumes that VLMs work as a proxy of human evaluation, and we demonstrate both qualitative and quantitative improvement of the video quality. It would be important to extend it to human feedback, not limited to VLMs. In practice, we often do implicit or explicit best-of-N sampling for video generation. In contrast, DLBS-LA exhibits much better computational efficiency. Spending more computation at inference time significantly improves perceptual quality, but it is orthogonal compared to speeding up the sampling process via distillation [61, 62], architecture changes [63], or parallel sampling [64]. We believe both high-quality and speedy sampling have practical needs and should be balanced.
@@ 8 Conclusion
This paper studies which metrics we should optimize and how to optimize them for better text-to-video generation. We point out that feedback from humans or capable VLMs reflects multiple dimensions of video quality, so optimizing an existing metric alone is insufficient; rather, we should calibrate the reward by combining. Our DLBS with LA estimator reduces the error propagation from the inaccurate reward estimate. We demonstrate DLBS is the most scalable, efficient and robust inference-time search that significantly improves video quality under the same computational costs. We hope our work encourages more uses of inference-time computation for text-to-video models.
@@ Acknowledgements
We thank Po-Hung Yeh, Shohei Taniguchi, Kuang-Huei Lee, Arnaud Doucet, Heiga Zen, and Yusuke Iwasawa for their support and helpful discussion on the initial idea of this work. HF was supported by JSPS KAKENHI Grant Number JP22J21582 (by March, 2025), and MS was supported by JSPS KAKENHI Grant Number JP23H04974.
Appendix A Broader Impacts
@@
On the other hand, the ability to generate highly realistic videos raises concerns about the potential for misuse in creating deceptive content, including deepfakes and misinformation. Like other generative models, such as large language models, text-to-video models, and their inference-time search, may inherit and amplify biases present in the training data due to the misalignment. This might lead to the generation of videos that perpetuate harmful stereotypes or underrepresent certain groups.
@@
Appendix B Implementation Details
B.1 Code.
Our implementation for the experiment is available at https://anonymous.4open.science/r/T2V-Diffusion-Search-537B.
B.2 Models.
Our experiments cover three text-to-video diffusion models.
  • Latte [18]: a T5-conditioned Latent Diffusion Transformer with 1.1 B parameters, built on PixArt-$\alpha$ [65, 66].
  • CogVideoX [19]: a larger DiT-based model with 2B or 5B parameters. We mainly used CogVideoX-5B.
  • Wan 2.1 [20]: a DiT-based flow model with 1.3B or 14B parameters. We use Wan 2.1-1.3B for reward calibration and Wan 2.1-14B for inference-time search experiments.
B.3 Hyperparameters
  • Latte: DDIM scheduler with a linear noise schedule $(\beta_{\text{start}}=1.0\times 10^{-4},\ \beta_{\text{end}}=2.0\times 10^{-2})$ and classifier-free guidance scale $w_{\text{cfg}}=7.5$.
  • CogVideoX: DDIM scheduler with the original settings and $w_{\text{cfg}}=6.0$. Owing to computational resource constraints, we limit the frame length to 17 per sample.
  • Wan 2.1: DPMSolver++ with guidance scale $w_{\text{cfg}}=5.0$. For the same computational reasons, the spatial resolution is limited to $832\times 480$ and the frame length to 33 per sample.
B.4 Hardware configuration
  • Latte: FP16 inference on a single NVIDIA A100 (40 GB), batch size 1.
  • CogVideoX: BF16 inference on a single NVIDIA A100 (40 GB), batch size 1.
  • Wan 2.1: FP16 inference on four NVIDIA H100s (80 GB each), batch size 1.
B.5 AI-feedback endpoints
We use API endpoints: gemini-1.5-pro-002 and gpt-4o-2024-11-20.
Appendix C Details of Metric Rewards
Subject Consistency  We adopt the subject consistency metric proposed in VBench [12] to quantify how consistently a subject is depicted across consecutive video frames. Concretely, for each frame $i$ in a video, we extract a feature representation $\mathbf{d}_{i}$ using DINO [46] with a ViT-B/16 [67] backbone. Let $\langle\mathbf{d}_{i},\mathbf{d}_{j}\rangle$ denote the cosine similarity between the features $\mathbf{d}_{i}$ and $\mathbf{d}_{j}$. Then, VBench defines the subject consistency metric as follows:
@@ Appendix C Details of Metric Rewards
Motion Smoothness  We adopt the frame-interpolation-based metric originally proposed in VBench [12] to assess whether a generated video’s motion is smooth and physically plausible. In particular, this metric leverages the motion prior from AMT [47], employing its AMT-S variant for frame reconstruction. Concretely, let $\bigl{[}\mathbf{f}_{0},\mathbf{f}_{1},\mathbf{f}_{2},\ldots,\mathbf{f}_{2n}\bigr{]}$ denote the frames of a generated video. We remove each odd-numbered frame to obtain a lower-frame-rate sequence $\bigl{[}\mathbf{f}_{0},\mathbf{f}_{2},\mathbf{f}_{4},\ldots,\mathbf{f}_{2n}\bigr{]}$, and rely on AMT-S to reconstruct the missing frames $\bigl{[}\hat{\mathbf{f}}_{1},\hat{\mathbf{f}}_{3},\ldots,\hat{\mathbf{f}}_{2n-1}\bigr{]}$. We then compute the Mean Absolute Error (MAE) between these reconstructed frames and the original odd-numbered frames, denoting this measure by $R_{\text{smoothness}}$. Finally, following the normalization scheme introduced in VBench, we define:
@@
Dynamic Degree  This measure quantifies the overall magnitude of dynamic object movement. Let $T$ be the total number of frames in the generated video. For each pair of consecutive frames $t$ and $t+1$, we estimate the optical flow $\mathbf{v}_{t}$ using RAFT [48], compute its norm $\|\mathbf{v}_{t}\|$, and sum these values across all frames:
@@
Aesthetic Quality  This criterion evaluates compositional rules, color harmony, and overall artistic merit on a per-frame basis. Concretely, for each frame $i$ in a video, we extract a CLIP image embedding $\mathbf{c}^{image}_{i}$ using the CLIP ViT-L/14 model [68]. We then feed $\mathbf{c}^{image}_{i}$ into the LAION aesthetic predictor [49], which assigns a raw rating $r_{i}\in[0,10]$. To normalize these scores to the $[0,1]$ range, we set
@@
Imaging Quality  This indicator assesses low-level distortions (e.g., over-exposure, noise, blur) in each generated frame. We adopt the MUSIQ predictor [50], trained on the SPAQ dataset [69]. The frame-wise score is normalized to [0, 1] by dividing by 100, and the final video score is the mean of these normalized values across all frames in the same way as Equation 12 and Equation 13.
Text-Video Consistency This measure captures how closely a generated video’s content aligns with its text prompt. We employ ViCLIP [51], a model pre-trained on a 10M video-text dataset and fine-tuned to handle temporal relationships, to embed both the video frames and the text. Since ViCLIP computes embeddings from 8-frame inputs, we sample 8 frames from each video. Let $\mathbf{v}^{\text{video}}$ denote the resulting video embedding and $\mathbf{v}^{\text{text}}$ denote the text embedding. We then define the final alignment score as the cosine similarity between these embeddings:
@@
Appendix D Algorithms with Continuous-time Diffusion Process
In this section, we present our algorithms for a continuous-time diffusion process. For Wan 2.1 [20], we integrated the proposed search algorithm into DPMSolver++ [70], a widely used continuous-time solver. Algorithms 3 and 4 present the pseudocode. Although we present the first-order variant for clarity, the procedure extends straightforwardly to higher-order formulations. Throughout this section, we adopt the notation of Lu et al. [70].
Algorithm 3 Diffusion Latent Beam Search (DLBS) with SDE-DPMSolver++
0:  signal prediction latent diffusion model $z_{\theta}$, reward function $r^{\prime}$, time steps $\{t_{s}\}_{s=0}^{M}$, noise scheduling parameter $\{\alpha_{t_{s}}\}_{s=0}^{M},\{\sigma_{t_{s}}\}_{s=0}^{M}$, number of beams $B$, number of candidates $K$
1:  $\textbf{z}^{1}_{t_{0}},\cdots,\textbf{z}^{B}_{t_{0}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$   $\triangleright$ Initial $B$ beams
2:  for $s=1$ to $M$ do
3:     for $j=1$ to $B$ do
4:         $\triangleright$ Update one step to produce $\mathbf{z}_{t_{s}}^{j}$
5:        $\textbf{z}^{j}_{t_{s}}=\frac{\sigma_{t_{s}}}{\sigma_{t_{s-1}}}e^{-h}\textbf{z}_{t_{s-1}}^{j}+\alpha_{t_{s}}(1-e^{-2h})z_{\theta}(\textbf{z}_{t_{s-1}}^{j})$
6:     end for
7:     if $s<M$ then
8:        for $j=1$ to $B$ do
9:            $\triangleright$ Sample $K$ next candidate latents
10:           $\mathbf{z}_{t_{s}}^{ij}=\mathbf{z}^{j}_{t_{s}}+\sigma_{t}\sqrt{e^{-2h}-1}\epsilon_{t_{s-1}}^{i}$ with $\epsilon_{t_{s-1}}^{1},...,\epsilon_{t_{s-1}}^{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
11:            $\triangleright$ Estimate the clean sample from noisy latent
12:           $\hat{\mathbf{z}}_{t_{M}|t_{s}}^{ij}=\frac{\sigma_{t_{M}}}{\sigma_{t_{s}}}\mathbf{z}_{t_{s}}^{ij}-\alpha_{t_{M}}(e^{-h}-1)z_{\theta}(\mathbf{z}^{ij}_{t_{s}}))$
13:        end for
14:         $\triangleright$ Search $B$ higher-reward beams from $KB$ latents
15:        $\texttt{budget}:=\{(\mathbf{z}_{t_{s}}^{11},\hat{\mathbf{z}}_{t_{M}|t_{s}}^{11}),\cdots,(\mathbf{z}_{t_{s}}^{KB},\hat{\mathbf{z}}_{t_{M}|t_{s}}^{KB})\}$
16:        for $j^{\prime}=1$ to $B$ do
17:           $\mathbf{z}_{t_{s}}^{j^{\prime}}=\operatorname{argmax}_{\mathbf{z}^{ij}_{t_{s}}\in\texttt{budget}}~{}r^{\prime}(\hat{\mathbf{z}}^{ij}_{t_{M}|t_{s}})$
18:           $\texttt{budget}=\texttt{budget}\setminus\{(\mathbf{z}_{t_{s}}^{j^{\prime}},\hat{\mathbf{z}}^{\operatorname{argmax}}_{t_{M}|t_{s}})\}$
19:        end for
20:        $j\in\{1,\cdots,B\}\leftarrow j^{\prime}$   $\triangleright$ Reset selected $B$ indices
21:     end if
22:  end for
23:  return: $\mathbf{z}_{t_{M}}=\operatorname{argmax}_{\mathbf{z}^{j}_{t_{M}}\in\{\textbf{z}^{1}_{t_{M}},\cdots,\textbf{z}^{B}_{t_{M}}\}}~{}r^{\prime}(\mathbf{z}^{j}_{t_{M}})$
Algorithm 4 Lookahead (LA) with DPMSolver++
0:  signal prediction latent diffusion model $z_{\theta}$, current diffusion latent $\textbf{z}_{t_{s}}$, number of lookahead steps $M^{\prime}(<<M)$
1:   $\triangleright$ Run $M^{\prime}$-step deterministic DPMSolver++ starting from $\textbf{z}_{t_{s}}$
2:  $\tilde{s}(u)\in\{s,\dots,\lfloor\frac{M^{\prime}-u}{M^{\prime}}s+\frac{u}{M^{\prime}}M\rfloor,\dots,\lfloor\frac{1}{M^{\prime}}s+\frac{M^{\prime}-1}{M^{\prime}}M\rfloor,M\}$
3:  Select new lookahead noise schedule $\{\tilde{\alpha}_{t_{u}}\}_{u=0}^{M^{\prime}}$ for $M^{\prime}$-step interpolation of the rest of original $\{\alpha_{t_{s}^{\prime}}\}_{s^{\prime}=0}^{M}$
4:  $\textbf{z}_{t_{\tilde{s}(0)}}:=\textbf{z}_{t_{s}}$
5:  for $u=1$ to $M^{\prime}$ do
6:     $\tilde{\textbf{z}}_{t_{\tilde{s}(u)}|t_{\tilde{s}(u-1)}}=\frac{\sigma_{t_{\tilde{s}(u)}}}{\sigma_{t_{\tilde{s}(u-1)}}}\mathbf{z}_{t_{\tilde{s}(u-1)}}^{ij}-\alpha_{t_{\tilde{s}(u)}}(e^{-h}-1)z_{\theta}(\mathbf{z}^{ij}_{t_{\tilde{s}(u-1)}}))$
7:  end for
8:  return: $(\textbf{z}_{t_{s}},\tilde{\textbf{z}}_{t_{M}|t_{\tilde{s}(0)}})$triangleright Latent and LA estimator
Appendix E Theoretical Analysis on Lookahead Estimator
Consider the Lookahead estimator described in Algorithm 2, which obtains the state $\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}$ by performing $T^{\prime}$ steps of DDIM (or another diffusion-based sampling) with $\eta=0.0$. Our goal is to show that, as $T^{\prime}$ grows, the reward estimate $r^{\prime}(\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)})$ converges to $r^{\prime}(\mathbf{z}_{0})$, thereby improving estimation accuracy.
Let $\mathbf{z}_{t-1}$ be a state in the latent space from which we wish to recover the initial latent $\mathbf{z}_{0}$. By applying $T^{\prime}$ steps of DDIM with $\eta=0.0$, we obtain an approximation $\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}$. From prior work [71], the error $\|\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}-\mathbf{z}_{0}\|$ scales as follows:
$\|\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}-\mathbf{z}_{0}\|\leq\begin{cases}\mathcal{O}(1/T^{\prime})&\text{(DDIM)},\\ \mathcal{O}(1/\sqrt{T^{\prime}})&\text{(DDPM)},\\ \mathcal{O}(1/(T^{\prime})^{n})&\text{(an $n$-th order solver)}.\end{cases}$
Hence, increasing $T^{\prime}$ yields a progressively better approximation of $\mathbf{z}_{0}$.
Assume $\mathbf{z}_{0}$ is the latent representation at time $t=0$. By the Continuous Mapping Theorem, if $\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}\to\mathbf{z}_{0}$ as $T^{\prime}\to\infty$, then for any continuous function $f$, we have
$f(\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)})\;\to\;f(\mathbf{z}_{0}).$
Setting $f(\cdot)=r^{\prime}(\cdot)$, where $r^{\prime}$ is our reward model, yields
$r^{\prime}(\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)})\;\to\;r^{\prime}(\mathbf{z}_{0}),$
as $T^{\prime}\to\infty$.
We further assume that the reward model $r^{\prime}(\cdot)$ is Lipschitz continuous with Lipschitz constant $L$. Then for any two latent states $\mathbf{z}_{a}$ and $\mathbf{z}_{b}$, the reward estimates satisfy
$|r^{\prime}(\mathbf{z}_{a})-r^{\prime}(\mathbf{z}_{b})|\;\leq\;L\|\mathbf{z}_{a}-\mathbf{z}_{b}\|.$
Hence, the order of the error in $r^{\prime}(\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)})$ tracks the order of the error in $\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}$ itself. Explicitly,
$\bigl{|}r^{\prime}(\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)})-r^{\prime}(\mathbf{z}_{0})\bigr{|}\;\leq\;L\,\|\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}-\mathbf{z}_{0}\|,$
implying that an $\mathcal{O}(1/T^{\prime})$ (or better) approximation in latent space implies an $\mathcal{O}(1/T^{\prime})$ (or correspondingly better) approximation in the reward space.
As $T^{\prime}$ increases, $\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)}$ converges to $\mathbf{z}_{0}$, and consequently $r^{\prime}(\tilde{\mathbf{z}}_{0\mid\tilde{t}(0)})$ converges to $r^{\prime}(\mathbf{z}_{0}$. Because the reward model is Lipschitz continuous, this convergence ensures that the error in reward estimation decreases at the same order as the error of the latent approximation. Therefore, employing the LA estimator with a larger $T^{\prime}$ yields a more accurate reward estimate.
Appendix F List of Prompts